OpenStudy (anonymous):

4 (x+7) = 6 (x - 5) solving proportions

OpenStudy (anonymous):

x=29

OpenStudy (stormfire1):

4x+28=6x-30 4x-6x=-30-28 -2x=-58 x=29

OpenStudy (anonymous):

thanks so much

OpenStudy (stormfire1):

np :)

OpenStudy (anonymous):

here is another one can i have a metle